OpenStudy (anonymous):

solve for x: 3+(x+3)^(1/2) =x help please?

OpenStudy (superdavesuper):

Squaring both left n right hand side to get rid of the ^(1/2) first...

OpenStudy (anonymous):

so that would make it 9+x+3=x^2?

OpenStudy (superdavesuper):

Almost but u miss one term on the left: (a+b)^2=a^2 + 2ab + b^2